OpenStudy (anonymous):

f(x)=4-(12/(x^2))Calculate the lower and upper sum estimates of the area under the curve over the interval [2,5] using subintervals of width ½.

ganeshie8 (ganeshie8):

what have you tried ?

OpenStudy (anonymous):

I found delta x

OpenStudy (anonymous):

I am not sure how to approach the problem though.

ganeshie8 (ganeshie8):

basically you need to estimate the area under given curve using rectangles for left sum you need to find area of below rectangles http://gyazo.com/7ae848ca5057eb28c8933bf534c1c481

OpenStudy (anonymous):

I know that you use six rectangles. Every video and website says you have to use Riemann sum, but that is in the next lesson.

OpenStudy (anonymous):

Do you have to use a calculator or is there and equation I can use?

ganeshie8 (ganeshie8):

you may use calculator notice that you just need to find the area of those 6 rectangles and add up

ganeshie8 (ganeshie8):

what did you get for \(\Delta x\) ?

OpenStudy (anonymous):

I don't now if deltax is 6 or 3/n

OpenStudy (anonymous):

.5

ganeshie8 (ganeshie8):

hey we're given \(\Delta x = \frac{1}{2}\)

OpenStudy (anonymous):

So delta x is the width?

ganeshie8 (ganeshie8):

yes, then total number of rectangles = \(\large \dfrac{5-2}{\frac{1}{2}} = ?\)

OpenStudy (anonymous):

See i tried doing this problem a harder way, that's why i am confused.

ganeshie8 (ganeshie8):

its okay :) can we say the area using left sum is below :\[\Delta x \left[ f(2) + f(2+\Delta x) + f(2 + 2\Delta x) + \cdots + f(2+5\Delta x)\right]\] ?

OpenStudy (anonymous):

yes?

ganeshie8 (ganeshie8):

replacing \(\Delta x\) by \(\frac{1}{2}\) gives \[\frac{1}{2} \left[ f(2) + f(2+\frac{1}{2} ) + f(2 + 2\cdot \frac{1}{2} ) + \cdots + f(2+5\cdot \frac{1}{2} )\right]\]

ganeshie8 (ganeshie8):

see if you can evaluate ^

OpenStudy (anonymous):

See i know that for the left sum you start at 2 and end with 4.5 and the right you use 3 and end with 5

ganeshie8 (ganeshie8):

fcor the right sum you start with 2.5

OpenStudy (anonymous):

Ya sorry i meant 2.5

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

since its intervals of .5

OpenStudy (anonymous):

do we just plug those numbers into the original equation now?

ganeshie8 (ganeshie8):

yep

OpenStudy (anonymous):

Alright I get thanks

OpenStudy (anonymous):

I have no idea what is was doing, i was using

ganeshie8 (ganeshie8):

you're welcome :) just ask if something doesnt make sense still right sum will give you overestimate : http://gyazo.com/86dafae0a550f6d06fbd86f606ecc9e2

OpenStudy (anonymous):

\[\sum_{i=1}^{n}\]

OpenStudy (anonymous):

When do you use what i just posted?

ganeshie8 (ganeshie8):

both are same

OpenStudy (anonymous):

I get the concept of how to do it its just the technical stuff that confuses me. So n is .5 then right?

ganeshie8 (ganeshie8):

Left sum : \[\begin{align} &\Delta x \left[ f(2) + f(2+\Delta x) + f(2 + 2\Delta x) + \cdots + f(2+5\Delta x)\right] \\~\\&= \sum\limits_{i=1}^6 f(2 + (i-1)\Delta x)\times \Delta x\\~\\&=\sum\limits_{i=1}^6 f(2 + (i-1)\frac{1}{2})\times \frac{1}{2}\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

n = number of rectangles = (5-2)/0.5 = 6

OpenStudy (anonymous):

That's what exactly what I was doing, so this is the same thing as we just did?

ganeshie8 (ganeshie8):

Yep

OpenStudy (anonymous):

This method is just so much more confusing.Any way thanks for your help i really appreciate it.

ganeshie8 (ganeshie8):

keep in mind you're just multiplying `width` and the `height` for area of each rectangle : \[\begin{align}& f(2 + (i-1)\Delta x)\times \Delta x\\~\\ \end{align}\] then adding them up : \[\begin{align}& \sum\limits_{i=1}^{6} f(2 + (i-1)\Delta x)\times \Delta x\\~\\ \end{align}\]

ganeshie8 (ganeshie8):

`riemann sum` is just a fancy name for above method ^

OpenStudy (anonymous):

I now know that, thanks.Also would the right sum be 3i/n +2

OpenStudy (anonymous):

That's why i am confused, Riemann sum is the next lesson. Any way thanks again.

ganeshie8 (ganeshie8):

right sum would be \[\begin{align}& \sum\limits_{i=1}^{6} f(2 + i\Delta x)\times \Delta x\\~\\ \end{align}\] how did u get 3i/n+2 ?

OpenStudy (anonymous):

Well my OS is broken, but